# Alignment and Adversarial Robustness: Are More Human-Like Models More Secure?

### v1

@@ Abstract
Representational alignment refers to the extent to which a model’s internal representations mirror biological vision, offering insights into both neural similarity and functional correspondence. Recently, some more aligned models have demonstrated higher resiliency to adversarial examples, raising the question of whether more human-aligned models are inherently more secure. In this work, we conduct a large-scale empirical analysis to systematically investigate the relationship between representational alignment and adversarial robustness. We evaluate 118 models spanning diverse architectures and training paradigms, measuring their neural and behavioral alignment and engineering task performance across 106 benchmarks as well as their adversarial robustness via AutoAttack. Our findings reveal that while average alignment and robustness exhibit a weak overall correlation, specific alignment benchmarks serve as strong predictors of adversarial robustness, particularly those that measure selectivity towards texture or shape. These results suggest that different forms of alignment play distinct roles in model robustness, motivating further investigation into how alignment-driven approaches can be leveraged to build more secure and perceptually-grounded vision models.
@@ 1 Introduction
A longstanding goal in computer vision is to develop models that process images in a way that aligns with human perception. Representational alignment—how closely a model resembles biological vision—has been studied extensively with the goal of measuring, bridging, or increasing alignment in machine learning models Sucholutsky et al. (2024). Recent observations suggest that alignment may have implications beyond neuroscience: models that are more aligned with human perception have also exhibited increased robustness to adversarial examples—inputs with near-imperceptible perturbations that induce model misclassification— Dapello et al. (2020); Li et al. (2019), hinting at a deeper connection between alignment and security.
@@
However, the relationship between representational alignment and adversarial robustness remains poorly understood. While the former seeks to align models with human cognition, adversarial examples in security highlight a fundamental misalignment: imperceptible perturbations can drastically degrade model accuracy while leaving human perception unaffected. Prior robustness techniques, such as adversarial training Madry et al. (2019), are computationally expensive and potentially vulnerable to new attack strategies. Meanwhile, alignment research has not systematically examined whether more human-aligned models are inherently more robust to adversarial attacks. A fundamental question remains: do these objectives complement each other, leading to better-aligned and more robust models, or do they introduce conflicting trade-offs?
@@
In this work, we investigate the relationship between human alignment and robustness to adversarial examples in vision models through a diverse, large-scale empirical analysis. In our analysis, we study 118 models across different architectures and training schemes, measure their alignment across 106 different benchmarks on neural, behavioral, and engineering tasks via the BrainScore library Schrimpf et al. (2018). We then evaluate the adversarial robustness of these models using AutoAttack Croce & Hein (2020), a state-of-the-art ensemble attack.
In analyzing the correlations between model robustness and alignment, our findings reveal that while robustness is weakly correlated with vision alignment on average, certain alignment benchmarks serve as strong indicators of model robustness. Specifically, we find that the top six benchmarks that were most positively correlated with robust accuracy, even with strong perturbations, all measured a model’s selectivity towards texture. These results suggest that different forms of alignment play distinct roles in model robustness, motivating further investigation into how alignment-driven approaches can be leveraged to build more secure and perceptually-grounded vision models.
@@ 2 Background
Representational Alignment. Representational alignment studies the extent to which internal representations of machine learning models correspond to human cognitive processes. Early studies found that deep neural networks (DNNs) trained on large-scale image datasets develop hierarchical feature representations similar to those observed in the primate ventral stream, particularly in high-level visual areas like the inferior temporal (IT) cortex Yamins et al. (2013); Schrimpf et al. (2018). This led to efforts to quantify the alignment between artificial and biological vision, using techniques such as Representational Similarity Analysis (RSA) Kriegeskorte et al. (2008) and Centered Kernel Alignment (CKA) Kornblith et al. (2019). Current research in the area primarily focuses on measuring, bridging, and increasing both neural and behavioral alignment.
To improve alignment, researchers have proposed strategies that incorporate cognitive constraints or psychological priors into model architectures Dapello et al. (2020). Supervised fine-tuning with human-annotated datasets Dosovitskiy et al. (2021) ensures that learned representations align more closely with human-understandable features. Furthermore, novel techniques Muttenthaler et al. (2023); Li et al. (2019); Cheng et al. (2024) have been developed to encourage similarity between model activations and human neural responses as recorded through fMRI and EEG experiments. In this study, we use a comprehensive set of neural, behavioral, and engineering alignment metrics to quantify representational alignment.
Adversarial Examples. Although machine learning models have shown strong capabilities in achieving high accuracy across various tasks Liu et al. (2022); Dosovitskiy et al. (2021); Krizhevsky et al. (2017); He et al. (2016), they remain vulnerable to adversarial examples Croce & Hein (2020); Madry et al. (2019); Carlini & Wagner (2017); Goodfellow et al. (2015); Sheatsley et al. (2023). Adversarial examples are specially crafted inputs that contain perturbations which are imperceptible to humans, yet significantly decrease model accuracy. In computer vision systems, there have been many studies on developing attack algorithms, such as FGSM Goodfellow et al. (2015), PGD Madry et al. (2019), and AutoAttack Croce & Hein (2020). These methods aim to maximize model’s loss subject to constraints of perturbations defined by certain $\ell_{p}$-norms as follows:
@@
where $x$ and $y$ represent the original image and its predicted label, respectively, $\delta$ is the perturbation to solve for, and $L$ is the model’s loss function. The perturbation constraint $\epsilon$ is measured through an $\ell_{p}$-norm—most commonly $\ell_{\infty}$. While many works have historically evaluated the robustness of their model through the PGD attack Madry et al. (2019), it has been shown that “robust” models can often suffer from gradient masking, causing gradient-based attacks like PGD to fail Athalye et al. (2018), and leading to a sense of overestimated robustness. To overcome this, multiple attacks, including both white- and black-box attacks should be used Carlini et al. (2019). Thus, the AutoAttack ensemble Croce & Hein (2020) has become the de-facto standard for evaluating robustness.
@@ 3 Methods
Alignment. To measure alignment and download candidate models, we leverage the BrainScore Schrimpf et al. (2018) library. BrainScore provides a standardized framework for evaluating model similarity to biological vision through a set of neural, behavioral, and engineering benchmarks, supplying 106 benchmarks in total. These benchmarks quantify how closely a model’s internal representations and outputs correspond to neurophysiological recordings, human psychophysical behavior, and performance on engineered vision tasks. Neural alignment is measured by comparing activations from DNNs to neural recordings from primate visual cortex regions (e.g., V1, V2, V4, and IT), using similarity metrics like Representational Similarity Analysis (RSA) Kriegeskorte et al. (2008). Behavioral alignment assesses whether models replicate human psychophysical responses in object recognition and perturbation tests, while engineering alignment evaluates model robustness to controlled distortions, such as contrast reductions, or performance on out of distribution data.
@@
In total, the BrainScore library has documented benchmark scores for 434 models. Out of those, there are 197 models available in their registry (the remaining 237 models were either submitted privately or have been deprecated). From the 197 models in the registry, we removed an additional 72 models because either loading the model produced a ClientError due to a moved or removed model hosting location or the model was incompatible with ImageNet (either does not output 1000 classes or expects video streams). After this, we had to discard an additional 7 models, which represented all the VOne class models Dapello et al. (2020) because they were not able to run on AutoAttack due to gradient alteration or masking, suggesting that previous results finding that VOne models are more robust to adversarial examples could have been due to overestimated robustness and highlighting the importance of evaluating robustness under comprehensive attack strategies. After this filtering process, we were left with 118 models (see Appendix A) for our evaluation.
@@
Robustness. To evaluate the robustness of our models, we use AutoAttack Croce & Hein (2020); Croce et al. (2021), which serves as the standard for evaluating the robustness of neural networks due to its strong attack performance and fully automated parameter-free design. AutoAttack contains 4 attacks: APGD-CE, APGD-DLR, FAB, and Square Attack. By evaluating on AutoAttack, we are not only evaluating on the most performant attacks, but also integrating in both white-box attacks and black-box attacks which has been recommended in previous works to combat reporting overestimated robustness due to gradient masking or obfuscation Carlini & Wagner (2017).
@@
To better understand how the relationship between adversarial robustness and alignment changes as attacks change, we evaluate the $\ell_{\infty}$ robustness of our models at three different epsilon levels: $\epsilon=\{\frac{0.25}{255},\frac{0.5}{255},\frac{1}{255}\}$ to represent adversaries at different capability levels and small, medium, and large image distortion levels. While these values are typically lower than what would be benchmarked on platforms such as RobustBench Croce et al. (2021), we choose these values with the goal of having a wide distribution of robust accuracies to identify separability between models, rather than the goal of bringing the model down to 0% accuracy as what is typically done.
@@ 4 Results
To facilitate our experiments, we use the BrainScore library v2.2.4 to measure alignment Schrimpf et al. (2018) and load models. Details on models evaluated can be found in Section 3. Once these models have been loaded and their alignment has been measured across the 106 alignment benchmarks, we evaluate their robustness using AutoAttack Croce & Hein (2020) from the TorchAttacks Kim (2021) library v3.5.1. The ImageNet Russakovsky et al. (2015) validation set is used for clean inputs to the model and serves as the starting point to generate adversarial examples. All experiments are run across 12 A100 GPUs with 40 GB of VRAM and CUDA version 11.1 or greater.
@@ 4.1 Average Alignment
We first investigate how well different classes of alignment predict the robustness of a model. Here, we study neural alignment, behavioral alignment, and engineering task performance. For each of these classes, we take the average score across all the benchmarks, giving us a single score for each model in the class. While many works have typically studied average vision alignment overall (i.e., the average of all the benchmarks across all classes), it has been shown that this can overemphasize behavioral alignment at the cost of neural alignment Ahlert et al. (2024). For each model, we then compute its robust accuracy against AutoAttack at 3 different values of epsilon $\epsilon=\{0.001,0.00196,0.00392\}$, which corresponds to $\{\frac{0.25}{255},\frac{0.5}{255},$ and $\frac{1}{255}\}$, respectively.
In Figure 1, we analyze the average score for neural alignment, behavioral alignment, and engineering task performance on the x-axis and the robust accuracy on the y-axis. Each dot represents a model, and the 3 colors correspond to the model’s robust accuracy at 3 different epsilon values. We compute the line-fit of the data at each epsilon value and report the statistical significance.
We find statistically significant correlations at: all $\epsilon$ values for neural alignment (explaining up to 13% of variance), $\epsilon=0.001$ for behavioral alignment (7.1% of variance), and at the two lowest $\epsilon$ values for engineering task performance (up to 27% of variance). Overall, the relatively low $R^{2}$ values, coupled with the difficulty of getting statistically significant correlations at higher epsilon values, suggests that average alignment scores are, at best, a weak indicator of robust accuracy.
@@ 4.2 Individual Benchmarks
Motivated by the previous experiment where we find that average alignment is weakly correlated with robust accuracy, we hypothesize this counter-intuitive result occurs because averaging scores across different benchmarks may obscure that some individual benchmarks are stronger predictors of robust accuracy than others. To further explore this hypothesis, we collect all models’ scores on individual benchmarks for the three classes (neural alignment, behavioral alignment, and engineering task performance) and compute the correlation between each of these scores and robust accuracy at our three different $\epsilon$ values. Figure 2 shows a heatmap of the 106 different benchmarks on the x-axis and robust accuracy at three different $\epsilon$ values on the y-axis. In each cell, we report the Pearson correlation coefficient between the selected benchmark score and robust accuracy across models.
Figure 2: Heatmap of each of the BrainScore benchmarks, ordered and separated (black bars) by area of alignment (neural, behavioral, engineering) vs the robust accuracy. Each cell represents the correlation between a benchmark across models and the robust accuracy for those models.
@@
From this figure, we find multiple interesting trends. First, we see a wide range of correlations between different benchmarks, confirming our hypothesis that not every current alignment metric is a good indicator of robust accuracy. Additionally, we sometimes see significant changes to the correlation of robust accuracy and a benchmark as the $\epsilon$ value increases (and thus becomes a stronger attack). These changes appear to cluster by class of alignment. Roughly speaking, the neural alignment benchmarks (shown from the first to second black bar) tend to have more stable (and more positive) correlations as $\epsilon$ increases. The behavioral benchmarks (shown from the second to third black bars) tend to be, surprisingly, often negatively correlated with robust accuracy at mid and high $\epsilon$ values, and the correlation mostly decreases as $\epsilon$ increases. Finally, engineering task performance (shown from the third black bar to the end of the figure) seems highly dependent on the task, with benchmarks in this category having correlations at both ends of the spectrum.
@@
Interestingly, we find some trends in the benchmarks that strongly correlate with robust accuracy. Most notably, many of the benchmarks that exhibit strong positive correlations with robust accuracy even at high values of epsilon tend to measure a models bias toward texture to some degree. In the neural category, we found strong postive correlations in FreemanZiemba2013.V1-pls and FreemanZiemba2013.V2-pls from Freeman et al. (2013), which measures neural responses in V1 and V2 to naturalistic texture stimuli. In the engineering category, two sets of benchmarks stood out as having strong correlations with robust accuracy. First is the Geirhos2021cueconflict-top1 benchmark from Geirhos et al. (2019), which measures the probability of a model classifying an object using shape information rather than texture via texture-shape conflicted images. The other is the set of benchmarks from Hermann et al. (2020): Hermann2020cueconflict-shape_bias and Hermann2020cueconflict-shape_match, which similarly measures the probability of a model classifying an object using shape information and the percentage of the times the model classifies according to the shape class, rather than texture or other classes.

### v2

@@ Abstract
A small but growing body of work has shown that machine learning models which better align with human vision have also exhibited higher robustness to adversarial examples, raising the question: can human-like perception make models more secure? If true generally, such mechanisms would offer new avenues toward robustness. In this work, we conduct a large-scale empirical analysis to systematically investigate the relationship between representational alignment and adversarial robustness. We evaluate 144 models spanning diverse architectures and training paradigms, measuring their neural and behavioral alignment and engineering task performance across 105 benchmarks as well as their adversarial robustness via AutoAttack. Our findings reveal that while average alignment and robustness exhibit a weak overall correlation, specific alignment benchmarks serve as strong predictors of adversarial robustness, particularly those that measure selectivity toward texture or shape. These results suggest that different forms of alignment play distinct roles in model robustness, motivating further investigation into how alignment-driven approaches can be leveraged to build more secure and perceptually-grounded vision models.
[…]
5395
@@ 1 Introduction
Representational alignment—how closely a model resembles biological vision—has been studied extensively with the goal of measuring, bridging, or increasing alignment in machine learning models [37]. Recent observations [10, 25] suggest that alignment may have implications beyond neuroscience: models that are more aligned with human perception have also exhibited increased robustness to adversarial examples—inputs with near-imperceptible perturbations that induce model misclassification— hinting at a deeper connection between alignment and security.
@@
However, the relationship between representational alignment and adversarial robustness remains poorly understood. While the former seeks to align models with human cognition, adversarial examples in security highlight a fundamental misalignment: imperceptible perturbations can drastically degrade model accuracy while leaving human perception unaffected. Prior robustness techniques, such as adversarial training [27], are computationally expensive and potentially vulnerable to new attack strategies. Although recent work has suggested links between perceptual alignment and robustness, especially in models incorporating biological priors [10, 25], a broad and systematic evaluation of this relationship across diverse models and alignment benchmarks is still lacking. A fundamental question remains: do these objectives complement each other, leading to better-aligned and more robust models, or do they introduce conflicting trade-offs?
@@
In this work, we investigate the relationship between human alignment and robustness to adversarial examples in vision models through a diverse, large-scale empirical analysis. In our analysis, we study 144 models across different architectures and training schemes, and measure their alignment across 105 different benchmarks on neural, behavioral, and engineering tasks via the BrainScore library [32]. We then evaluate the robustness of these models using AutoAttack [8], a state-of-the-art ensemble attack.
Our analysis reveals that while robustness is only weakly correlated with average vision alignment, specific alignment benchmarks are strongly predictive of adversarial robustness. We also find that models with similar alignment profiles exhibit similar robustness. Together, these findings suggest that different forms of alignment contribute differently to robustness, highlighting the value of alignment-driven approaches for improving security in vision systems.
  • Average alignment, particularly for behavioral benchmarks, poorly predicts robustness (explaining 6% of variance), demonstrating that not all methods of alignment increase robustness.
  • Individual benchmarks are much better indicators of robust accuracy, but the specific benchmark matters; we found instances of benchmarks that contributed positively and ones that contributed negatively to robustness in every alignment category.
  • The top benchmarks that are most indicative of robust accuracy show clear trends that inform the factors leading to higher robustness: (1) robust models process texture information specifically more similarly to humans than non-robust models to and (2) models’ ability to recognize objects without global structures intact actually hurts their robustness.
  • t-SNE visualizations reveal that robustness is structured in alignment space, with similar models forming robustness-consistent clusters and demonstrating that alignment on certain benchmarks is a good indicator of robustness.
In summary, we uncover specific features of alignment that closely tie with model robustness and show that increasing alignment on these benchmarks offers a new avenue for building more robust and human-aligned models.
@@ 2 Background
2.1 Representational Alignment
Representational alignment studies the extent to which internal representations of machine learning models correspond to human cognitive processes. Early studies found that deep neural networks (DNNs) trained on large-scale image datasets develop hierarchical feature representations similar to those observed in the primate ventral stream, particularly in high-level visual areas like the inferior temporal (IT) cortex [39, 32]. This led to efforts to quantify the alignment between artificial and biological vision, using techniques such as Representational Similarity Analysis (RSA) [23] and Centered Kernel Alignment (CKA) [22]. To improve alignment, researchers have proposed strategies that incorporate cognitive constraints or psychological priors into model architectures [10]. Supervised fine-tuning with human-annotated datasets [11] ensures that learned representations align more closely with human-understandable features. Furthermore, novel techniques [28, 25, 7] have been developed to encourage similarity between model activations and human neural responses as recorded through fMRI and EEG experiments. In this study, we use a comprehensive set of neural, behavioral, and engineering alignment metrics to quantify representational alignment.
2.2 Adversarial Examples
Although machine learning models have shown strong capabilities to achieve high accuracy in various tasks [26, 11, 24, 17], they remain vulnerable to adversarial examples [8, 27, 5, 16, 33]. Adversarial examples are specially crafted inputs that contain perturbations which are imperceptible to humans, yet can significantly decrease model accuracy. In computer vision systems, there have been many studies on developing attack algorithms, such as FGSM [16], PGD [27], and AutoAttack [8]. These methods aim to maximize model’s loss subject to constraints of perturbations defined by certain $\ell_{p}$-norms:
@@ 2.2 Adversarial Examples
where $x$ and $y$ represent the original image and its predicted label, respectively, $\delta$ is the perturbation to solve for, and $L$ is the model’s loss function. The perturbation constraint $\epsilon$ is measured through an $\ell_{p}$-norm—most commonly $\ell_{\infty}$. While many works have historically evaluated the robustness of their model through PGD [27], it has been shown that “robust” models can often suffer from gradient masking, causing gradient-based attacks like PGD to fail [2], and leading to a sense of overestimated robustness. To overcome this, multiple attacks, including both white- and black-box attacks should be used [6]. Thus, the AutoAttack ensemble [8] has become the de-facto standard for evaluating robustness.
@@ 3 Methods
This section details how we measure the alignment and robustness of machine learning models with the goal of assessing if more human-like vision models are also more resilient to security vulnerabilities.
3.1 Alignment.
@@
To measure alignment and download candidate models, we leverage the BrainScore [32] library. BrainScore provides a standardized framework for evaluating model similarity to biological vision through a set of neural, behavioral, and engineering benchmarks, supplying 106 benchmarks in total. These benchmarks quantify how closely a model’s internal representations and outputs correspond to neurophysiological recordings, human psychophysical behavior, and performance on engineered vision tasks. Neural alignment is measured by comparing activations from DNNs to neural recordings from primate visual cortex regions (e.g., V1, V2, V4, and IT), using similarity metrics like Representational Similarity Analysis (RSA) [23]. Behavioral alignment assesses whether models replicate human psycho-physical responses in object recognition and perturbation tests, while engineering alignment evaluates model robustness to controlled distortions, such as contrast reductions, or performance on out-of-distribution data. We use 105 benchmarks by discarding one of them, which has NaN value for all the selected models.
@@
In total, the BrainScore library has documented benchmark scores for 434 models. Out of those, there are 240 models available in their registry (the remaining models were either submitted privately or have been deprecated). From the 240 models in the registry, we further removed an additional 89 models that were either incompatible with ImageNet (e.g., do not output 1000 classes or expect video streams) or could not be run on RobustBench due to gradient masking. After this, we had to discard an additional 7 models, which represented all the VOne class models [10] because they were not able to run on AutoAttack due to gradient alteration or masking, suggesting that previous results finding that VOne models are more robust to adversarial examples could have been due to overestimated robustness and highlighting the importance of evaluating robustness under comprehensive attack strategies. After this filtering process, we were left with 144 models for our evaluation.
Model diversity is critical for evaluating the generalization of alignment-robustness relationships. Thus, the 144 models selected for our evaluation are representative by covering a broad spectrum of architectures and training recipes. The majority of architectures are convolutional neural networks (CNNs) [29] such as various ResNet [17] and VGG variants [34]. We also include more recent architectural designs such as Vision Transformers (ViTs) [11], which uses self-attention mechanisms to capture global dependencies in images, its variants [4], and hybrid models [26, 38] (i.e., a combination of CNNs and ViTs). This architectural diversity is complemented by models trained with different strategies such as standard supervised learning and self-supervised learning (e.g., contrastive learning [30]). Notably, the evaluated models also include those specifically designed or trained with properties relevant to the study’s central hypothesis: some models have undergone adversarial training [27], which improves robustness against adversarial examples, while others use mechanisms to emphasize shape bias [13, 14, 18], shifting model reliance from texture to shape information. This design has been shown to be more aligned with human perception. The comprehensive model set allows for a thorough exploration of how architectural design and training methods impact the relationship between representational alignment and adversarial robustness.
3.2 Robustness.
@@
To evaluate the robustness of our models, we use AutoAttack [8, 9], which serves as the standard for evaluating the robustness of neural networks due to its strong attack performance and fully automated parameter-free design. AutoAttack contains 4 attacks: APGD-CE, APGD-DLR, FAB, and Square Attack. By evaluating on AutoAttack, we are not only evaluating on the most performant attacks, but also integrating in both white-box attacks and black-box attacks which has been recommended in previous works to combat reporting overestimated robustness due to gradient masking or obfuscation [5].
@@
To better understand how the relationship between adversarial robustness and alignment changes as attacks change, we evaluate the $\ell_{\infty}$ robustness of our models at three different epsilon levels: $\epsilon=\{\frac{0.25}{255},\frac{0.5}{255},\frac{1}{255}\}$ to represent adversaries at different capability levels and small, medium, and large image distortion levels. While these values are typically lower than what would be benchmarked on platforms such as RobustBench [9], we choose these values with the goal of having a wide distribution of robust accuracies to identify separability between models, rather than the goal of bringing the model down to 0% accuracy as is typically done.
@@ 4 Results
Figure 2: Heatmap of each of the BrainScore benchmarks, ordered and separated (black bars) by area of alignment (neural, behavioral, engineering) vs the robust accuracy. Each cell represents the correlation between a benchmark and the robust accuracy across models.
To address this, we conduct a series of experiments. We use the BrainScore library v2.2.4 to measure alignment [32] and load models. Details on models evaluated can be found in Section 3. Once these models have been loaded and their alignment has been measured across the 105 alignment benchmarks, we evaluate their robustness using AutoAttack [8] from the TorchAttacks [21] library v3.5.1. The ImageNet [31] validation set is used for clean inputs to the model and serves as the starting point to generate adversarial examples. All experiments are run across 12 A100 GPUs with 40 GB of VRAM and CUDA version 11.1 or greater. Our code is available for download at https://github.com/kyangl/alignment-security.
@@ 4.1 Average Alignment
We first investigate how well different classes of alignment—neural alignment, behavioral alignment, and engineering task performance—predict model robustness. For each class, we take the average score across all the benchmarks, giving us a single score for each model in the class. While many works have typically studied average vision alignment overall (i.e., the average of all the benchmarks across all classes), it has been shown that this can overemphasize behavioral alignment at the cost of neural alignment [1].
We then assessed each model’s robust accuracy against AutoAttack at three different values of epsilon $\epsilon=\{0.001,0.00196,0.00392\}$, which corresponds to $\{\frac{0.25}{255},\frac{0.5}{255},\frac{1}{255}\}$, respectively. In Figure 1, we analyze the average score for neural alignment, behavioral alignment, and engineering task performance on the x-axis and the robust accuracy on the y-axis. Each dot represents a model, and the 3 colors correspond to the model’s robust accuracy at the three different epsilon values. We perform linear fits for each epsilon value and report the statistical significance.
We find statistically significant ($p\leq 0.05$) correlations for: the two lowest $\epsilon$ values for neural alignment (explaining up to 9% of variance), $\epsilon=0.001$ for behavioral alignment (6% of variance), and at the two lowest $\epsilon$ values for engineering task performance (up to 19% of variance). Overall, the relatively low $R^{2}$ values, coupled with the difficulty of getting statistically significant correlations at higher epsilon values, suggest that average alignment scores are, at best, a weak indicator of robust accuracy. This counter-intuitive finding leads us to study further on individual benchmarks.
@@ 4.2 Individual Benchmarks
Motivated by the previous experiment where we find that average alignment is weakly correlated with robust accuracy, we hypothesize that averaging scores across different benchmarks may obscure that some individual benchmarks are stronger predictors of robust accuracy than others. Here, certain alignment benchmarks could be more indicative of robust accuracy than others.
@@
To examine this hypothesis, we collect all models’ scores on individual benchmarks for the three classes (neural alignment, behavioral alignment, and engineering task performance) and compute the correlation between each of these scores and robust accuracy at our three different $\epsilon$ values. Figure 2 shows a heatmap of the 105 different benchmarks on the x-axis and robust accuracy at three different $\epsilon$ values on the y-axis. In each cell, we report the Spearman correlation coefficient between the selected benchmark score and robust accuracy across models.
@@
From this figure, we find multiple interesting trends. First, we see a wide range of correlations between different benchmarks, confirming our hypothesis that not every current alignment metric is a good indicator of robust accuracy. Additionally, we sometimes see significant changes to the correlation of robust accuracy and a benchmark as the $\epsilon$ value increases (and thus becomes a stronger attack). These changes appear to cluster by class of alignment. Roughly speaking, the neural alignment benchmarks (shown from the first to the second black bar) seem highly dependent on the task, with benchmarks in this category having correlations at both ends of the spectrum. The behavioral benchmarks (shown from the second to third black bars) tend to be, surprisingly, often negatively correlated with robust accuracy at mid and high $\epsilon$ values, and the correlation mostly decreases as $\epsilon$ increases. Finally, engineering task performance (shown from the third black bar to the end of the figure) tends to have more stable (and more positive) correlations as $\epsilon$ increases.
Furthermore, we identified specific benchmarks that are strongly correlated with robust accuracy. Interestingly, many of these benchmarks that exhibit strong positive correlations with robust accuracy even at high epsilon values measure, to some degree, a model’s selectivity towards texture information, meaning that models which are more robust to adversarial examples tend to process texture information more similarly to humans. Below, we highlight and discuss the benchmarks we found to be most strongly (positive or negative) correlated with robust accuracy.
In the neural category, we found the correlation between robust accuracy and alignment to be highly dependent on the area of visual processing. At small to medium values of $\epsilon$, correlations are strongest in benchmarks that measure alignment to V4 and IT areas of the visual system. At later values of $\epsilon$, which represents a stronger attack, FreemanZiemba2013.V2-pls which measures responses to naturalistic texture stimuli in V2 [12], was the metric with the strongest positive correlation. This suggests that processing texture more similarly to biological vision systems in later visual areas is even more important than early visual areas to yield robust models. Interestingly, we found that metrics utilizing fMRI data were typically the most negatively correlated.
In the behavioral category, we see many strong positive correlations at low values of $\epsilon$. The sets of benchmarks that stand out the most are the BMD and Baker [3] groups, which measure the ability to recognize objects by silhouettes when the shape has been distorted. In Baker and Elder [3], it was found that humans are less able to recognize objects when shape information is distorted than neural networks are, demonstrating that machine learning models have a certain degree of insensitivity to changes in shape information and thus don’t rely as heavily on it as humans do for object recognition. This finding further supports the notion that neural networks may rely more on texture information rather than shape information [13]. The strong positive relationship between these benchmarks and performance on adversarial examples suggests that models’ ability to recognize objects without spatial relationships actually hurts their overall robustness.
@@
Finally, on engineering tasks, we see strong positive correlations across nearly all benchmarks with low $\epsilon$. However, these correlations decrease as $\epsilon$ increases for most benchmarks. The benchmarks that remain most strongly positively correlated with high $\epsilon$ are both related to bias towards shape over texture information. First is the Geirhos2021cueconflict-top1 benchmark [13], which measures the probability of a model classifying an object using shape information rather than texture via texture-shape conflicted images. The other is the set of benchmarks from [18]: Hermann 2020 cue conflict shape _bias and Hermann2020cueconflictshape_match, which similarly measures the probability of a model classifying an object using shape information and the percentage of the times the model classifies according to the shape class, rather than texture or other classes. In all, these results show that when models are able to classify inputs according to their shape information, as humans do, rather than texture information, they will be more robust to adversarial examples.
Across all categories of visual alignment, we find that models which exhibit a stronger reliance on shape information and robust processing of texture cues (i.e., more aligned to how humans process information) tend to be more resilient to adversarial examples. These results suggest that increasing alignment towards preferring and processing high-level visual features, particularly textures, in the way biological vision systems do, serves as a fruitful direction for creating more robust and more aligned models.
4.3 Robustness in Alignment Space
Building on our finding that certain alignment benchmarks (especially those measuring texture sensitivity) are strongly correlated with adversarial robustness, we next explore whether models with similar alignment profiles also exhibit similar levels of robust accuracy. The goal of this analysis is to see if model similarity on alignment benchmarks is predictive of model performance against adversarial examples, which would support the hypothesis that aligning models under certain metrics results in better robustness.
To compare the similarity of model performance across benchmarks, we utilize t-SNE to reduce the dimensions of the results and project them into a subspace that can be visualized. In Figure 3, each point represents one of our 144 models, colored by its robust accuracy under $\ell_{\infty}$ adversarial perturbations with $\epsilon=0.001$. We observe that the models are distributed across the embedding space with clustering patterns that are unique to their robustness.
Analyzing the non-robust models (blue dots), we find that while some cluster in the upper left, many are scattered broadly throughout the space. This dispersion suggests that poor robustness is not associated with a single alignment profile. Rather, there are many different ways in which models can fail to defend against adversarial examples. That is, vulnerability appears to be distributed across a range of misaligned configurations.
In contrast, the robust models (red dots) form a tight, well-separated cluster. These models not only share a similar alignment profile, but their separation from the rest of the models suggests that robust behavior is tied to a specific region of alignment space. This strongly supports the view that certain alignment characteristics are predictive of robustness.
These findings offer a new perspective on adversarial robustness; rather than unique quirks of the model resulting in vulnerability to adversarial examples, there are specific, isolated properties that allow models to more robustly process inputs. In other words, it’s not that specific misalignment will lead to model vulnerability, but rather specific alignment will lead to robustness. Furthermore, these properties can be measured in alignment metrics, and thus optimized for in order to build more robust models.
Figure 3: t-SNE plot of principal components from benchmark features for each model colored by the model’s robust accuracy.